### v1

# Neural Configuration-Space Barriers for Manipulation Planning and Control

Abstract
Planning and control for high-dimensional robot manipulators in cluttered, dynamic environments require both computational efficiency and robust safety guarantees. Inspired by recent advances in learning configuration-space distance functions (CDFs) as robot body representations, we propose a unified framework for motion planning and control that formulates safety constraints as CDF barriers. A CDF barrier approximates the local free configuration space, substantially reducing the number of collision-checking operations during motion planning. However, learning a CDF barrier with a neural network and relying on online sensor observations introduce uncertainties that must be considered during control synthesis. To address this, we develop a distributionally robust CDF barrier formulation for control that explicitly accounts for modeling errors and sensor noise without assuming a known underlying distribution. Simulations and hardware experiments on a 6-DoF xArm manipulator show that our neural CDF barrier formulation enables efficient planning and robust real-time safe control in cluttered and dynamic environments, relying only on onboard point-cloud observations.
I Introduction
Efficient planning and safe control for high-degree-of-freedom (DoF) robotic manipulators are a central challenge in robotics, particularly when operating in cluttered and dynamic environments with noisy sensor measurements. Sampling-based motion planners, such as rapidly-exploring random trees (RRT) [25, 42] and probabilistic roadmaps (PRM) [21], have gained widespread adoption for manipulator motion planning because they effectively explore high-dimensional spaces without exhaustive configuration-space discretization. Despite their strengths, these methods rely on frequent collision checks, making the efficiency and accuracy of collision checking a critical bottleneck. Moreover, in highly dynamic settings with sensor noise and rapidly changing obstacles, planning alone is insufficient to ensure safety during execution, promoting the integration of robust real-time control strategies.
To address these challenges, signed distance functions (SDFs) have gained popularity for encoding the distance from any point in the workspace to the boundary of an object, thereby enabling precise collision checks and facilitating gradient-based optimization. Recently, neural networks have been leveraged to learn SDF representations of both environments [35, 39, 30] and robot bodies [23, 28, 46], providing rich shape modeling along with efficient distance and gradient queries. Building on these advancements, neural SDFs have been adopted in motion planning [46, 36] and optimization-based safe control [30, 38], yielding promising results in complex, dynamic scenarios.
(a) Workspace visualization
(b) Bubble-CDF planning
[…]
While SDFs have proven effective in workspace-centric tasks, such as planning and obstacle avoidance for mobile robots [34, 30], recent work [27] demonstrates that directly representing distances in configuration space can be particularly advantageous for manipulators with high-dimensional kinematic chains. In a configuration distance field (CDF), the distance is measured in joint space between a given configuration and a workspace point, eliminating the need for repeated transformations between task and configuration spaces. Like neural SDFs, neural CDFs offer a compact and differentiable representation, supporting efficient gradient-based computations for real-time motion planning [27] and reactive control [7]. However, learning-based representations inevitably introduce modeling errors, and real-world deployments must also contend with sensor noise and other uncertainties.
Motivated by these challenges, we propose an approach that integrates a neural CDF barrier into both motion planning and control, addressing the twin requirements of computational efficiency and safety under uncertainty. The contributions of this work are summarized as follows.
  • We develop a motion planning algorithm that uses a neural CDF barrier to construct a graph of fully traversable spheres in the robot’s configuration space, allowing significantly reduced collision checks.
  • We purpose a distributionally robust control barrier function formulation from the neural CDF barrier to account for dynamic obstacles, CDF modeling inaccuracies, and sensor noise.
  • We validate our CDF-based planning and control algorithms in extensive simulations and real-robot experiments on a 6-DoF xArm manipulator in dynamic and cluttered environments.
II Related Work
This section reviews existing approaches for motion planning and control of robotic manipulators.
To plan and execute manipulator motions, the prevailing approach is to use discrete sampling-based planners such as RRT [25], PRM [21], and their variants [24, 19] to compute a global path in the configuration space, which is then followed using feedback controllers such as PID [20]. These methods have been widely adopted due to their simplicity and ability to generate collision-free paths in complex environments. However, as manipulators are deployed in increasingly dynamic and unstructured environments, the limitations of sampling-based planners become evident, particularly their reliance on computationally expensive collision checking, which restricts frequent replanning, and their inability to adapt paths in real-time, leading to potential safety risks during execution.
To address the challenges associated with frequent collision checking, a line of work focuses on constructing explicit approximations of the free configuration space. Methods of Dai et al. [9], Werner et al. [43], Deits and Tedrake [11] decompose the free configuration space into multiple convex polyhedra by solving a series of optimization problems, typically assuming that workspace obstacles are represented by geometric primitives such as spheres or cylinders. Some prior works, such as Ademovic and Lacevic [1], Danielson [10], utilize workspace distances to construct safe regions; however, these methods are restricted to simple manipulator geometries, such as infinitely thin lines. Additionally, spherical safe regions have been explored for efficient collision checking [5, 37], biased sampling in free workspace representations [6], and navigation function design [44]. Our approach differs by using these regions as convex constraints for trajectory optimization in configuration space, extending the concept introduced by Lee et al. [26] to high-dimensional robot configurations and more complex environments.
Accurate geometric modeling is essential for enabling manipulators to safely navigate complex and dynamic environments. Signed distance functions (SDFs) have been extensively utilized in manipulator motion planning and control due to their efficient computation of distances and gradients [23, 13, 46, 28]. Koptev et al. [23] introduced neural joint SDFs to facilitate collision-free motion planning for manipulators, while Vasilopoulos et al. [41] leveraged manipulator SDFs for reactive motion planning by integrating a model predictive path integral trajectory generator with a vector field-based follower. From a safety-critical perspective, Singletary et al. [38] proposed a CBF-based trajectory filtering strategy that dynamically adjusts pre-planned manipulator trajectories to account for updated collision constraints. Yu et al. [45] developed a CBF-induced neural controller (CBF-INC) that integrates with RRT for motion planning, utilizing state-based or point-cloud inputs to enhance success rates and reduce exploration nodes.
Recent advancements have extended the utility of distance fields to configuration space for manipulators. Li et al. [27] introduced the concept of a configuration distance field (CDF) and proposed using a neural network to approximate it, allowing the estimation of the joint space distance between a given configuration and a workspace point. Building on this, Chi et al. [7] leveraged neural CDFs as control barrier functions (CBFs) for safe control, demonstrating that CDF-based methods outperform SDF-based approaches due to their more reliable gradient estimates.
While Chi et al. [7] is closely related to this work, we differ in several key aspects. First, we extend the application of neural CDFs to both planning and control frameworks, addressing the dual objectives of planning efficiency and safety under uncertainty. Specifically, our planning strategy leverages configuration-space bubbles constructed with neural CDF barriers, enabling efficient exploration and trajectory optimization. Second, our control formulation explicitly accounts for the modeling errors inherent in neural approximations and sensor measurement uncertainties by incorporating a distributionally robust formulation. These enhancements enable our framework to provide robust and reliable performance in dynamic and cluttered environments while relying on point-cloud sensory measurements.
III Notation
The sets of real, non-negative real, and natural numbers are denoted by $\mathbb{R}$, $\mathbb{R}_{\geq 0}$, and $\mathbb{N}$, respectively. For $N\in\mathbb{N}$, we let $[N]:=\{1,2,\dots N\}$. The interior and boundary of a set $\mathcal{C}\subset\mathbb{R}^{n}$ are denoted by $\text{Int}(\mathcal{C})$ and $\partial\mathcal{C}$. The gradient of a differentiable function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is denoted by $\nabla V$. A continuous function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ is of extended class $\mathcal{K}_{\infty}$ if it is strictly increasing, $\alpha(0)=0$, and $\lim_{r\rightarrow\infty}\alpha(r)=\infty$. The special orthogonal group of dimension $p$ is denoted by $\text{SO}(p)=\{\mathbf{R}\in\mathbb{R}^{p\times p}\mid\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}_{p},\det(\mathbf{R})=1\}$. The special Euclidean group of dimension $p$, denoted $\text{SE}(p)$ is the set of matrices $\begin{pmatrix}\mathbf{R}&\mathbf{t}\\ \mathbf{0}^{\top}&1\end{pmatrix}$ with $\mathbf{R}\in\text{SO}(p)$ and $\mathbf{t}\in\mathbb{R}^{p}$.
IV Problem Formulation
We consider a robot manipulator with $m$ revolute joints and $m+1$ links. The joint configuration is represented by $\mathbf{q}=[q_{1},\ldots,q_{m}]^{\top}\in\mathcal{Q}$, where $\mathcal{Q}=\{\mathbf{q}\in\mathbb{R}^{m}\mid\mathbf{q}_{\min}\leq\mathbf{q}\leq\mathbf{q}_{\max}\}$ is the configuration space bounded by the joint limits. The forward kinematics of the manipulator map a joint configuration $\mathbf{q}$ to the end-effector pose in task space, represented as $\mathbf{T}_{\text{ee}}(\mathbf{q})\in\text{SE}(3)$.
Given a configuration $\mathbf{q}$, we denote the robot shape by a set-valued function $\mathcal{S}(\mathbf{q})\subset\mathbb{R}^{3}$, and its surface by $\partial\mathcal{S}(\mathbf{q})$. The kinematics of the manipulator are described by:
$\dot{\mathbf{q}}(t)=\mathbf{u}(t),\quad\mathbf{q}(0)=\mathbf{q}_{0},$ (1)
where $\mathbf{u}\in\mathcal{U}\subset\mathbb{R}^{m}$ is the joint velocity control input, and $\mathbf{q}_{0}$ is the initial configuration of the robot.
We consider a dynamic environment containing both static and dynamic obstacles. At time $t$, let $\mathcal{O}(t)\subset\mathbb{R}^{3}$ denote the closed set of obstacles, and let the free space be represented as $\mathcal{F}(t)=\mathbb{R}^{3}\setminus\mathcal{O}(t)$, which is an open set. We assume that a sensor (e.g., depth camera or LiDAR) provides point-cloud measurements $\mathcal{P}(t)\subset\mathbb{R}^{3}$, which approximate the boundary of the time-varying obstacle region $\mathcal{O}(t)$. The extrinsic parameters relating the sensor pose to the robot base are assumed to be calibrated offline.
Problem 1.
Consider a robot manipulator with initial configuration $\mathbf{q}_{0}$ and dynamics in (1). Given point-cloud observations $\mathcal{P}(t)$ of the obstacles $\mathcal{O}(t)$ in the environment, design a control policy that drives the robot’s end-effector $\mathbf{T}_{\text{ee}}$ to a desired goal pose $\mathbf{T}_{\text{G}}\in\text{SE}(3)$, while ensuring the robot body remains in free space, i.e., $\mathcal{S}(\mathbf{q}(t))\subset\mathcal{F}(t)$, $\forall t\geq 0$.
V Safe Barrier from Configuration-Space Distance Function
Efficient motion planning and control for robotic manipulators require accurate and scalable safety representations in configuration space. In this section, we first review the notion of a signed distance function (SDF) in the workspace, then introduce the CDF and its key properties, and finally define a neural CDF barrier for planning and control applications.
V-A Configuration Space Distance Function
Motivated by successful applications in safe mobile robot navigation [17, 30, 22], SDFs have recently been used for motion planning and control of robot manipulators [23, 28, 45, 31, 32]. An SDF $f_{s}$ measures the distance from a workspace point $\mathbf{p}\in\mathbb{R}^{3}$ to the robot surface $\partial\mathcal{S}(\mathbf{q})$:
$f_{s}(\mathbf{p},\mathbf{q})=\begin{cases}-\min\limits_{\mathbf{p}^{*}\in\partial\mathcal{S}(\mathbf{q})}\|\mathbf{p}-\mathbf{p}^{*}\|,&\text{if }\mathbf{p}\in\mathcal{S}(\mathbf{q}),\\ \phantom{-}\min\limits_{\mathbf{p}^{*}\in\partial\mathcal{S}(\mathbf{q})}\|\mathbf{p}-\mathbf{p}^{*}\|,&\text{if }\mathbf{p}\notin\mathcal{S}(\mathbf{q}).\end{cases}$ (2)
When represented using a differentiable model (e.g., neural network [35, 23]), this function enables efficient motion planning and collision avoidance by facilitating rapid computation of distances and gradients. While SDFs are widely applied to mobile robots to enforce workspace safety constraints, manipulators with high-dimensional joint spaces benefit from a configuration-centric perspective.
Li et al. [27] introduced a configuration distance function (CDF), which encodes the minimum joint-space distance (in radians) needed for a robot at configuration $\mathbf{q}$ to make contact with a point $\mathbf{p}$. Formally, given a robot SDF $f_{s}$, the robot CDF can be computed as
$f_{c}(\mathbf{p},\mathbf{q})=\min_{\mathbf{q}^{\prime}}\|\mathbf{q}-\mathbf{q}^{\prime}\|,\;\;\text{subject to}\;\;f_{s}(\mathbf{p},\mathbf{q}^{\prime})=0.$ (3)
Similar to an SDF, this CDF representation satisfies an Eikonal equation with respect to $\mathbf{q}$:
$\|\nabla_{\mathbf{q}}f_{c}(\mathbf{p},\mathbf{q})\|=1,$ (4)
whenever it is differentiable in $\mathbf{q}$.
By operating directly in joint space, CDF-based methods avoid iterative conversions between task and configuration spaces, and provide gradient information that is inherently meaningful for optimization-based joint control synthesis [7].
V-B Neural CDF Barrier
Building on the concept of the CDF, we define a CDF barrier to ensure the robot remains in a collision-free region of the configuration space. Formally, a CDF barrier associated with robot CDF $f_{c}$ and obstacle set $\mathcal{O}(t)$ is defined as:
$h(\mathbf{q},t):=\inf_{\mathbf{p}\in\partial\mathcal{O}(t)}f_{c}(\mathbf{p},\mathbf{q}),$ (5)
where $\partial\mathcal{O}(t)$ is the obstacle set boundary. Then, the time-varying safe set in configuration space induced by $h(\mathbf{q},t)$ is
$\mathcal{C}_{\text{safe}}(t)=\{\mathbf{q}\in\mathcal{Q}\mid h(\mathbf{q},t)\geq 0\}.$ (6)
In real-world settings, $\partial\mathcal{O}(t)$ must be inferred from point-cloud measurements $\mathcal{P}(t)$ from a depth camera or LiDAR. Because the points in $\mathcal{P}(t)$ are noisy samples from the boundary of $\mathcal{O}(t)$, we approximate the CDF barrier $h(\mathbf{q},t)$ as the minimum of $f_{c}(\mathbf{p},\mathbf{q})$ over the points in $\mathcal{P}(t)$:
$h(\mathbf{q},t)\;\approx\;\min_{\mathbf{p}\,\in\,\mathcal{P}(t)}f_{c}(\mathbf{p},\mathbf{q}).$ (7)
Moreover, computing $f_{c}(\mathbf{p},\mathbf{q})$ exactly can be difficult for high-DoF manipulators because it involves (infinitely) many potential contact configurations $\mathbf{q}^{\prime}$ in (3). Following Li et al. [27], we model $f_{c}$ by a multi-layer perceptron $\hat{f}_{c}(\mathbf{p},\mathbf{q};\boldsymbol{\theta})$ with learnable parameters $\boldsymbol{\theta}$. This representation reduces the storage requirements compared to volumetric or tabular encodings and supports parallel queries for a set of points $\mathbf{p}$ or configurations $\mathbf{q}$ with associated gradient computations. However, learning-based CDF representations inevitably introduce modeling errors. Combining the point-cloud approximation with the neural CDF approximation yields a practical CDF barrier:
$\hat{h}(\mathbf{q},t;\boldsymbol{\theta})\;=\;\min_{\mathbf{p}\,\in\,\mathcal{P}(t)}\hat{f}_{c}(\mathbf{p},\mathbf{q};\,\boldsymbol{\theta}),$ (8)
where the noise in $\mathcal{P}(t)$, its lack of complete coverage of the obstacle boundary, and the approximation error in $\hat{f}_{c}(\cdot;\boldsymbol{\theta})$ constitute multiple sources of uncertainty.
In the following sections, we leverage the fast parallel query capabilities of the neural CDF barrier to enable efficient sampling-based planning in configuration space (Sec. VI), and then propose a distributionally robust control strategy that tracks the planned trajectory while ensuring real-time safety despite uncertainties and dynamic obstacles (Sec. VII).
VI Configuration Space Bubble-Based Planning
Motion planning for high-DoF manipulators typically involves numerous collision checks [19], making global motion planning computationally expensive. To address this, we propose the bubble-CDF planner, a sampling-based approach that efficiently explores the configuration space while ensuring safety via neural CDF barriers. Unlike conventional RRT-based planners that rely on dense edge validation, our approach constructs configuration space bubbles as local safe regions, enabling rapid exploration and collision-free path generation.
In Problem 1, the robot must move from an initial configuration $\mathbf{q}_{0}$ to any valid joint configuration $\mathbf{q}_{\text{G}}$ that achieves a desired end-effector pose $\mathbf{T}_{\text{ee}}(\mathbf{q}_{\text{G}})=\mathbf{T}_{\text{G}}\in\mathrm{SE}(3)$. Inverse kinematics (IK) provides $K$ goal configurations $\{\mathbf{q}_{\text{G}}^{i}\}_{i\in[K]}$, each of which achieves the desired end-effector pose.
Our approach builds on the Rapidly Exploring Bubble Graph (RBG) algorithm [26], which constructs a configuration-space roadmap where each node represents a locally safe region (bubble) instead of a single configuration. The learned CDF barrier $\hat{h}(\mathbf{q},t)$ is used to define the bubble radius dynamically, ensuring rapid, collision-free exploration without incremental edge validation. This substantially reduces the collision-checking overhead in high-dimensional spaces while preserving the efficiency of sampling-based planners.
VI-A Rapidly-exploring Bubble Graph (RBG)
Suppose $\mathbf{q}\in\mathcal{Q}$ is a safe configuration. We aim to identify a radius $r(\mathbf{q})\geq 0$ such that any configuration $\mathbf{q}^{\prime}$ satisfying $\|\mathbf{q}-\mathbf{q}^{\prime}\|\leq r(\mathbf{q})$ is safe. We call this spherical region a configuration-space bubble and denote it as $\mathcal{B}(\mathbf{q},r(\mathbf{q}))$. Given the neural CDF barrier $\hat{h}(\mathbf{q},t)$ for fixed time $t$, we construct the bubble radius at configuration $\mathbf{q}$ as:
$r(\mathbf{q})=\hat{h}(\mathbf{q},t)-\eta,$ (9)
where $\eta>0$ is a safety margin. All configurations $\mathbf{q}^{\prime}$ within this radius satisfy the neural CDF barrier with margin $h(\mathbf{q}^{\prime},t)\geq\eta$ [26].
To explore a high-dimensional configuration space efficiently using these local safety certificates, RBG adapts the RRT algorithm [25] so that each vertex in the graph corresponds to a bubble rather than a single configuration. In RBG, the step size, corresponding to the distance the planned tree advances toward a sampled point, is dynamically set by the radius of each bubble. This obviates manual tuning of a fixed step size and reduces collision-checking overhead.
Algorithm 1 details how RBG constructs a bubble-based graph. RBG iteratively samples random configurations (biased toward goal configurations), identifies the nearest bubble in configuration space, and if the sample lies outside that bubble, creates a new bubble by querying the neural CDF barrier to determine its radius. The newly created bubble is then connected to any existing bubbles $\mathcal{B}(\mathbf{q}_{i},r(\mathbf{q}_{i}))$ with overlapping safe regions, i.e., if
$\|\mathbf{q}_{\mathrm{new}}-\mathbf{q}_{i}\|\;\leq\;r(\mathbf{q}_{\mathrm{new}})+r(\mathbf{q}_{i}).$ (10)
This process continues until the maximum number of bubbles $N_{\mathrm{max}}$ is reached or feasible paths to one goal configuration are identified.
Algorithm 1 Rapidly Exploring Bubble Graph
1:Start configuration $\mathbf{q}_{0}$, goal configurations $\{\mathbf{q}_{\mathrm{G}}^{i}\}_{i=1}^{K}$, Neural CDF barrier $\hat{h}(\mathbf{q},t)$, safety margin $\eta$, Max. no. of bubbles $N_{\mathrm{max}}$, Min. radius $r_{\mathrm{min}}$.
2:Graph $\mathcal{G}$ containing bubbles from start to goals
3:function BuildBubbleGraph($\mathbf{q}_{0},\{\mathbf{q}_{\mathrm{G}}^{i}\}$)
4:    $\mathcal{G}.V\leftarrow\{\mathcal{B}(\mathbf{q}_{0},\hat{h}(\mathbf{q}_{0},t)-\eta)\},\quad\mathcal{G}.E\leftarrow\emptyset$
5:    while $|\mathcal{G}.V|<N_{\mathrm{max}}$ do
6:         $\mathbf{q}_{\mathrm{rand}}\leftarrow\textsc{SampleRandom}()$triangleright With goal bias
7:         $\mathcal{B}_{\mathrm{near}}\leftarrow\textsc{NearestBubble}(\mathcal{G},\mathbf{q}_{\mathrm{rand}})$
8:         if $\mathbf{q}_{\mathrm{rand}}$ outside $\mathcal{B}_{\mathrm{near}}$ then
9:             $\mathbf{q}_{\mathrm{new}}\leftarrow\textsc{ExtendToward}(\mathcal{B}_{\mathrm{near}},\mathbf{q}_{\mathrm{rand}})$
10:             $r_{\mathrm{new}}\leftarrow\hat{h}(\mathbf{q}_{\mathrm{new}},t)-\eta$
11:             if $r_{\mathrm{new}}>r_{\mathrm{min}}$ then
12:                 $\mathcal{B}_{\mathrm{new}}\leftarrow\mathcal{B}(\mathbf{q}_{\mathrm{new}},r_{\mathrm{new}})$
13:                 UpdateConnections($\mathcal{G},\mathcal{B}_{\mathrm{new}}$)
14:    return $\mathcal{G}$
15:function UpdateConnections($\mathcal{G},\mathcal{B}_{\mathrm{new}}$)
16:    for $\mathcal{B}_{i}\in\mathcal{G}.V$ do
17:         if $\|\mathcal{B}_{\mathrm{new}}.\mathbf{q}-\mathcal{B}_{i}.\mathbf{q}\|\leq r(\mathbf{q}_{\mathrm{new}})+r(\mathbf{q}_{i})$ then
18:             $\mathcal{G}.E\leftarrow\mathcal{G}.E\,\cup\,(\mathcal{B}_{\mathrm{new}},\mathcal{B}_{i})$
19:    $\mathcal{G}.V\leftarrow\mathcal{G}.V\,\cup\,\{\mathcal{B}_{\mathrm{new}}\}$
VI-B Path Selection and Trajectory Optimization
Once the bubble-based graph $\mathcal{G}$ is constructed, we must extract and refine a collision-free path connecting the start configuration $\mathbf{q}_{0}$ to one of the $K$ goal configurations $\{\mathbf{q}_{\mathrm{G}}^{i}\}$. Since $\mathcal{G}$ may contain multiple feasible routes, we first select an optimal path according to a desired criterion (e.g., minimum total edge distance). We then solve a local trajectory optimization problem, fitting a continuous curve through the sequence of bubbles to ensure a smooth, dynamically feasible trajectory that remains collision-free.
Path Selection. We assign a cost to each edge in $\mathcal{G}$, for instance, the single-sided Hausdorff distance between two overlapping bubbles [26]. A shortest-path search (e.g., Dijkstra [12] or A* [18]) is run from the start bubble to every goal bubble. The path with the lowest overall cost is then passed to the optimization step, as summarized in Algorithm 2.
Trajectory Optimization. To obtain a smooth trajectory, we optimize a sequence of Bézier curves connecting the start configuration $\mathbf{q}_{0}$ to the selected goal configuration $\mathbf{q}_{\text{G}}^{\star}$ through the path of bubbles $\{\mathcal{B}_{1},\ldots,\mathcal{B}_{n}\}$. Each Bézier curve segment of degree $d$ is defined by its control points:
$\gamma_{j}(t)=\sum_{l=0}^{d}\binom{d}{l}t^{l}(1-t)^{d-l}\mathbf{c}_{j,l},$ (11)
where $\mathbf{c}_{j,l}\in\mathbb{R}^{m}$ is the $l$-th control point of the $j$-th curve segment. The $k$-th derivative of a Bézier curve is also a Bézier curve, and its squared norm integral can be expressed as:
[…]
where $\mathbf{c}_{j}$ is the vectorized form of all control points in segment $j$, and $\mathbf{Q}_{k}$ is a positive semidefinite matrix. This leads to the following convex quadratically constrained quadratic program:
[…]
subject to $\|\mathbf{c}_{j,l}-\mathbf{q}_{j}\|^{2}\leq r(\mathbf{q}_{j})^{2},\quad\forall j,l$
$\gamma_{j}(1)=\gamma_{j+1}(0),\quad j=1,\ldots,n-1$
$\gamma_{j}^{(k)}(1)=\gamma_{j+1}^{(k)}(0),\quad j=1,\ldots,n-1,\;k=1,2$
$\gamma_{1}(0)=\mathbf{q}_{0},\quad\gamma_{n}(1)=\mathbf{q}_{\text{G}}^{*},$
$\gamma_{1}^{(i)}(0)=\mathbf{0},\quad\gamma_{n}^{(i)}(1)=\mathbf{0},\quad i=1,\ldots,d$
where $\mathbf{c}_{j,i}$ are the control points of the $j$-th Bézier curve segment $\gamma_{j}$ of degree $d$, and $\gamma_{j}^{(k)}$ denotes its $k$-th derivative. The objective minimizes a weighted sum of integrated squared derivatives, promoting smoothness. The first constraint ensures safety by keeping control points within their respective bubbles, which by the convex hull property of Bézier curves guarantees the entire trajectory remains collision-free. The remaining constraints enforce continuity of position and derivatives between segments, as well as boundary conditions including zero velocity at endpoints. Unlike traditional sampling-based planners that may not have safety guarantees or require additional collision checks for post-optimized trajectory, our formulation provides a convex optimization problem that simultaneously guarantees smoothness and safety without requiring additional collision checking.
Algorithm 2 Path Selection and Trajectory Optimization
1:
2:Graph $\mathcal{G}$ with start bubble $\mathcal{B}_{\mathrm{start}}$, goal bubbles $\{\mathcal{B}_{\mathrm{goal}}^{i}\}_{i}$
3:Edge cost function $\textsc{EdgeCost}(\mathcal{B}_{i},\mathcal{B}_{j})$
4:Weights $\{w_{k}\}$ for trajectory cost
5:Smooth collision-free Bézier trajectory from $\mathbf{q}_{0}$ to some $\mathbf{q}_{\mathrm{G}}^{i}$
6:function PathSelectAndOptimize($\mathcal{G},\mathcal{B}_{\mathrm{start}},\{\mathcal{B}_{\mathrm{goal}}^{i}\}$)
7:    $\textit{bestCost}\leftarrow\infty,\quad\textit{bestPath}\leftarrow\textit{None},\quad i_{\mathrm{best}}\leftarrow-1$
8:    for $i=1\to K$ do
9:         $\textit{path}_{i}\leftarrow\textsc{ShortestPath}(\mathcal{G},\mathcal{B}_{\mathrm{start}},\mathcal{B}_{\mathrm{goal}}^{i},\textsc{EdgeCost})$
10:         if $\textit{path}_{i}$ exists then
11:             $\textit{cost}_{i}\leftarrow\textsc{PathCost}(\textit{path}_{i},\textsc{EdgeCost})$
12:             if $\textit{cost}_{i}<\textit{bestCost}$ then
13:                 $\textit{bestCost}\leftarrow\textit{cost}_{i}$, $\textit{bestPath}\leftarrow\textit{path}_{i}$, $i_{\mathrm{best}}\leftarrow i$
[…]
15:         return failure $\triangleright$ No feasible path found
16:    $\textbf{curves}\leftarrow\textsc{BezierOpti}(\textit{bestPath},\{w_{k}\})$triangleright Eq. (13)
17:    return $\bigl{(}\textbf{curves},\,i_{\mathrm{best}}\bigr{)}$
Throughout this section, we have introduced the bubble-CDF planner, which leverages neural CDF barriers to efficiently explore the configuration space while ensuring safety. By constructing configuration-space bubbles around safe configurations and dynamically adjusting step sizes based on bubble radii, our approach significantly reduces collision-checking overhead. Additionally, the trajectory optimization formulation in (13) guarantees smooth, dynamically feasible paths without requiring post-processing for collision checks.
Thus far, we have focused on global planning in static environments, assuming a time-invariant and accurate neural CDF barrier $\hat{h}(\mathbf{q},t)$. However, real-world deployments must contend with dynamic obstacles, CDF modeling errors, and sensor uncertainties. In Section VII, we introduce a distributionally robust control formulation to track the planned path in real time, explicitly accounting for these uncertainties to ensure safe and adaptive execution.
VII Distributionally Robust Manipulation Control
In this section, we present a control synthesis approach that guarantees safe real-time configuration trajectory tracking using imperfect neural CDF barrier estimates (due to sensing and modeling errors) even in the presence of dynamic obstacles. We first describe our approach to track the reference trajectory (Sec. VII-A). We then incorporate a distributionally robust control barrier function (DR-CBF) [33] to ensure safe control synthesis despite dynamic obstacles and neural CDF barrier uncertainty (Sec. VII-B).
VII-A Governor-based Trajectory Tracking
We first present our approach for tracking the planned piecewise Bézier configuration trajectory $\gamma:[0,1]\to\mathcal{Q}$, obtained in Sec. VI-B, in the absence of safety constraints. Inspired by Garone and Nicotra [16], Li et al. [29], we design a reference governor to determine a dynamically changing reference configuration $\gamma(s)$ along the trajectory, where $s\in[0,1]$ is a progress variable. In other word, $\gamma(0)=\mathbf{q}_{0}$ is the initial robot configuration, while $\gamma(1)=\mathbf{q}_{\text{G}}$ is one of the goal configurations. The governor system adapts the value of the progress variable $s$ based on the manipulator’s current configuration to ensure smooth progress along the trajectory. We then stabilize the manipulator configuration $\mathbf{q}$ to $\gamma(s)$ using a proportional-derivative (PD) controller.
The dynamics of $s$ are designed to balance the robot’s tracking capabilities with smooth progression along the trajectory:
$\dot{s}=\frac{k}{1+\|\mathbf{q}-\gamma(s)\|}(1-s^{\zeta}),$ (14)
where $k>0$ is a gain parameter, $\zeta\geq 1$ controls the rate of saturation as $s$ approaches 1, and $\|\mathbf{q}-\gamma(s)\|$ is the Euclidean distance between the current and reference configuration. This adaptive mechanism accelerates $s$ when the manipulator is near the reference and decelerates it when farther away, ensuring smooth progression without abrupt jumps.
To have the manipulator track the reference trajectory $\gamma$, we employ a PD controller as the nominal controller, which stabilizes the manipulator to the current local goal $\gamma(s)$:
$\bar{\mathbf{u}}(\mathbf{q},t)=-K_{P}(\mathbf{q}-\gamma(s))-K_{D}\dot{\mathbf{q}},$ (15)
where $K_{P}$ and $K_{D}$ are positive definite matrices controlling the proportional and derivative contributions, respectively. The first term ensures that the robot is drawn towards $\gamma(s)$, while the second term provides damping to minimize oscillations.
VII-B Distributionally Robust Neural CDF Barrier
To ensure safe operation while tracking the planned trajectory, the robot’s configuration $\mathbf{q}$ must remain within the safe set $\mathcal{C}_{\text{safe}}(t)$ as defined in (6). We utilize the CDF barrier as control barrier functions (CBFs) [3] by considering its time derivative. This leads to the following time-varying control barrier condition (CBC),
$\text{CBC}(\mathbf{q},\mathbf{u},t)$ $\!:=\!\nabla_{\mathbf{q}}h(\mathbf{q},t)^{\top}\mathbf{u}\!+\!\frac{\partial h(\mathbf{q},t)}{\partial t}\!+\!\alpha_{h}(h(\mathbf{q},t))$ (16)
$=\underbrace{[\nabla_{\mathbf{q}}h(\mathbf{q},t)^{\top},\;\;\alpha_{h}(h(\mathbf{q},t)),\;\;\frac{\partial h(\mathbf{x},t)}{\partial t}]}_{\boldsymbol{\xi}^{\top}(\mathbf{q},t)}\begin{bmatrix}\mathbf{u}\\ 1\\ 1\end{bmatrix},$
where $\alpha_{h}$ is a class $\mathcal{K}_{\infty}$ function.
With the nominal PD control law $\bar{\mathbf{u}}(\mathbf{q},t)$ in (15), one can enforce safety by solving CBF quadratic program (QP) [2] that generates the safe manipulator velocity control:
$\mathbf{u}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m}}\|\mathbf{u}-\bar{\mathbf{u}}(\mathbf{q},t)\|^{2}$ (17)
s.t. $\;\text{CBC}(\mathbf{q},\mathbf{u},t)\geq 0.$
However, the CDF barrier $h(\mathbf{q},t)$ and its gradients are not known exactly but rather are approximated through the neural CDF barrier $\hat{h}(\mathbf{q},t)$ in (8). Due to uncertainties introduced by the neural network model and the point-cloud data, $\hat{h}$ cannot be treated as an exact representation of $h$. Instead, we work with samples of the associated uncertainty vector $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$, defined in (16), which capture the variations in $\hat{h}$, its gradient, and its time derivative. The practical approach of obtaining such samples is discussed in Sec. VII-C.
Drawing inspiration from recent advances in distributionally robust optimization (DRO) [14, 33], we reformulate the CBC in a distributionally robust manner:
$\inf_{\mathbb{P}\in\mathcal{M}_{N}^{r}}\mathbb{P}\left(\text{CBC}(\mathbf{q},\mathbf{u},\boldsymbol{\xi})\geq 0\right)\geq 1-\epsilon,$ (18)
where $\mathcal{M}_{N}^{r}$ is a Wasserstein ambiguity set constructed from $\{\boldsymbol{\xi}_{i}\}_{i=1}^{N}$ and $\epsilon$ is a user-specified risk tolerance. To formally define $\mathcal{M}_{N}^{r}$, let $\Xi\subseteq\mathbb{R}^{m+2}$ denote the space of uncertainty vectors $\boldsymbol{\xi}$, and let $\mathcal{P}_{p}(\Xi)$ be the space of all Borel probability measures over $\Xi$ with finite $p$-th moment, by utilizing the $p$-Wasserstein distance $W_{p}$ [14], the ambiguity set
$\mathcal{M}_{N}^{r}:=\{\mu\in\mathcal{P}_{p}(\Xi)\;|\;W_{p}(\mu,\mathbb{P}_{N})\leq r\}$ (19)
defines a ball of distributions around the discrete empirical distribution $\mathbb{P}_{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{\boldsymbol{\xi}_{i}}$ of the samples.
This DRO formulation provides probabilistic safety guarantees by ensuring the safety constraint holds with high probability for all distributions around $\mathbb{P}_{N}$. The Wasserstein ambiguity set measures the distribution similarity and its parameters (radius $r$ and risk tolerance $\epsilon$) offer flexibility to account for different levels of uncertainty due to sensor noise and neural model inaccuracies, without requiring explicit or precise knowledge of the true distribution of $\boldsymbol{\xi}$.
Next, by utilizing the $1$-Wasserstein distance and following a convex reformulation of the DR-CBF in [33, Proposition 6.5], lead to the following optimization problem, whose feasibility implies the DRO constraint in (18):
$\mathbf{u}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m},s\in\mathbb{R},\beta_{i}\in\mathbb{R}}\|\mathbf{u}-\bar{\mathbf{u}}(\mathbf{q},t)\|^{2},$ (20)
s.t. $\;r\|\mathbf{u}\|_{\infty}\leq s\epsilon-\frac{1}{N}\sum_{i=1}^{N}\beta_{i},$
[…]
The reformulation maintains the linear dependence of the constraints on $\mathbf{u}$, ensuring that the optimization remains a QP and can be solved efficiently in real time. By integrating the DR-CBF constraint into the control synthesis, the proposed DR-CBF-QP safety filter enables robust and safe operation in dynamic environments.
VII-C Uncertainty Sample Selection
This section details our approach for obtaining samples $\{\boldsymbol{\xi}_{i}\}_{i=1}^{N}$ of the uncertainty vector $\boldsymbol{\xi}$ in the CBC in (16).
The definition of the neural CDF barrier $\hat{h}(\mathbf{q},t)$ in (8) depends on the point-cloud data $\mathcal{P}(t)$. The point cloud is represented as $\mathcal{P}(t)=\{\mathbf{p}_{j}\}_{j=1}^{M_{1}}$, where each $\mathbf{p}_{j}\in\mathbb{R}^{3}$ corresponds to a workspace point captured at time $t$. The velocity of each point, denoted as $\frac{\partial\mathbf{p}_{j}}{\partial t}$, is also considered to account for dynamics in the environment.
To account for uncertainty in the neural CDF model $\hat{f}(\mathbf{p},\mathbf{q})$, we employ Monte Carlo (MC) dropout in the network architecture [15]. This allows us to perform $M_{2}$ stochastic forward passes through the robot CDF neural network, collecting multiple realizations of $\hat{f}_{c}(\mathbf{p},\mathbf{q};\boldsymbol{\theta})$, along with its gradients. Specifically, for each forward pass of $f_{c}^{(i)}$, we evaluate $f_{c}^{(i)}$ across all $M_{1}$ points in $\mathcal{P}(t)$. Consequently, we obtain $M_{1}\times M_{2}$ realizations of:
$\left\{f_{c}^{(i)}(\mathbf{p}_{j},\mathbf{q}),\nabla_{\mathbf{q}}f_{c}^{i}(\mathbf{p}_{j},\mathbf{q}),\nabla_{\mathbf{p}}f_{c}^{(i)}(\mathbf{p},\mathbf{q})^{\top}\frac{\partial\mathbf{p}_{j}}{\partial t}\right\},$ (21)
where $i\in[M_{2}]$ corresponds to the MC dropout realization and $j\in[M_{1}]$ corresponds to the point in $\mathcal{P}(t)$.
From the $M_{1}\times M_{2}$ realizations, we select $N\ll M_{1}\times M_{2}$ samples in the set (21) based on the $N$ minimal values of
$\alpha_{h}(f_{c}^{(i)}(\mathbf{p}_{j},\mathbf{q}))+\nabla_{\mathbf{p}}f_{c}^{(i)}(\mathbf{p},\mathbf{q})^{\top}\frac{\partial\mathbf{p}_{j}}{\partial t},$ (22)
which effectively identifies samples $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$.
In this section, we introduced a DR-CBF control formulation that ensures safe manipulation under uncertainty. Unlike naive safety margin adjustments, our approach explicitly accounts for uncertainties in the neural CDF model, its gradient estimates, and point-cloud velocity observations. By leveraging a Wasserstein ambiguity set, we optimize for the worst-case uncertainty within a data-driven confidence region, providing stronger safety guarantees than standard CBF formulations. This enables adaptive and robust control in dynamic environments without requiring explicit probabilistic models of uncertainty.
VIII Evaluation
This section evaluates our planning and control techniques, which utilize neural CDF barriers, in simulations on 2-DoF planar arm and 6-DoF xArm as well as in real-world experiments on 6-DoF xArm. We evaluate the computational efficiency and solution quality of the bubble-CDF planner and the robustness of the DR-CBF controller in ensuring safety under dynamic obstacles and uncertainty.
VIII-A Baselines and Parameters
We compare our approach against the following baselines.
  • Planners:
    • CDF-RRT: RRT with collision checks based on the neural CDF barrier.
    • SDF-RRT [46]: RRT with collision checks using a neural SDF barrier.
    • SDF-RRT-Connect [24]: A variant of SDF-RRT with a bi-directional search strategy.
    • SDF-Lazy-RRT [19]: A variant of SDF-RRT with reduced collision-checking overhead.
  • Controllers:
    • PD Controller: A proportional-derivative controller $\bar{\mathbf{u}}(\mathbf{q},t)$ focused solely on trajectory tracking.
    • PD + CBF-QP [7]: A PD controller $\bar{\mathbf{u}}(\mathbf{q},t)$ augmented with a CBF-QP safety filter, where the CBF is constructed using the neural CDF barrier. However, this approach disregards uncertainty in the CDF model or sensor measurements, unlike ours.
The baseline planners are implemented using the Open Motion Planning Library (OMPL) [40], ensuring consistency across all algorithms. All controllers are implemented using Casadi [4], and the CBF-QP and our DR-CBF-QP problems are solved using the interior point optimizer.
Evaluation Metrics: For planners, we evaluate the number of collision checks and path length. For controllers, the evaluation metrics include success rate (percentage of trials where the robot safely reaches the goal) and tracking error (measured as the Fréchet distance between planned and executed paths).
Parameters: Tables I and II summarize the parameters used for the planners and controllers.
TABLE I: Planner parameters used for evaluation. Goal bias specifies the probability of sampling configurations directly toward the goal. Step size defines the maximum extension distance during tree expansion. Safety margin indicates the minimum clearance required for collision checking, and collision check resolution determines the granularity of edge validity checks, expressed as a fraction of the configuration space diagonal. For the 2-DoF planar robot, a resolution of 0.01 corresponds to 0.089 radians, while for the 6-DoF xArm, a resolution of 0.002 corresponds to 0.04 radians.
Planner Goal Bias Step Size Safety Margin Col. Check Resolution
Baselines 0.1 0.1 0.05 0.01 (2-DoF) / 0.002 (6-DoF)
Bubble-CDF 0.1 N/A 0.05 N/A
TABLE II: Control parameters used for evaluation. The PD control gains $K_{P}$ and $K_{D}$ are positive definite matrices, assumed to be diagonal with identical values along the diagonal. The function $\alpha_{h}$ is a linear class $\mathcal{K}_{\infty}$ function used in the CBF formulation. Parameters $k$ and $\zeta$ define the governor dynamics in (14). The risk tolerance $\epsilon$ and sample size $N$ correspond to the DR-CBF in (18), while $r$ represents the Wasserstein radius.
Parameter $K_{P}$ $K_{D}$ $\alpha_{h}$ $k$ $\zeta$ $\epsilon$ $N$ $r$
Value 0.8 0.1 1.0 0.2 12 0.1 10 0.02
VIII-B 2-DoF Planar Robot Simulation
(a) Initial Configuration
(b) Defensive Maneuver
(c) Resuming Tracking
(d) Goal Reached
Figure 2: Snapshots of a 2-link arm navigating a dynamic environment with purple obstacles (velocity directions shown by arrows). The arm is tasked to follow the planned path in Fig. 1. The arm is shown in blue and the trajectory of its end-effector is shown in red, and the local reference configuration goal $\gamma(s)$ is shown in green.
TABLE III: Planning performance comparison on a 2-DoF planar robot.
Planner Col. Checks Path Length
CDF-RRT $2108.9\pm 438.7$ $3.64\pm 0.69$
SDF-RRT [46] $2118.3\pm 388.9$ $3.61\pm 0.69$
SDF-RRT-Connect $2144.6\pm 459.7$ $3.63\pm 0.74$
SDF-Lazy-RRT $1759.2\pm 299.3$ $3.75\pm 0.83$
Bubble-CDF $\boldsymbol{153.8\pm 62.2}$ $3.72\pm 0.81$
We begin by evaluating our bubble-CDF planner against baseline sampling-based global motion planning approaches on a 2-DoF planar robotic arm. The arm consists of two links, each 2 meters long, with a fixed base located at the origin $(0,0)$. The joint angles are constrained within the range $[-\pi,\pi)$, and the initial configuration is set to $(0,0)$, placing the end-effector at $(4,0)$, as illustrated in Fig. 1.
Setup: We assess the performance of the bubble-CDF planner and all baselines in 500 randomly generated environments. Each environment includes $4$ static obstacles of varying sizes and positions, and the end-effector goal is randomly sampled to be at least $4$ meters away from its initial position while ensuring reachability. The robot receives point-cloud observations of the obstacle surfaces. Due to the 2-DoF planar structure, most reachable end-effector positions correspond to two distinct goal configurations, and the planners are tasked with finding a feasible path from the initial configuration to either one of the two goal configurations.
Results and Discussion: Table III summarizes the performance metrics across all 500 environments. The results demonstrate the significant efficiency of the bubble-CDF planner in reducing the number of collision checks while maintaining comparable path quality. The bubble-CDF planner requires $153.8\pm 62.2$ collision checks, while even the best baseline (Lazy-RRT) requires at least 10×10$\times$10 × more collision checks. This showcases the ability of our proposed bubble-CDF planner to explore the configuration space more efficiently. Despite this reduction, the bubble-based planner produces paths with similar lengths ($3.72\pm 0.81$) radians to those of the baselines.
TABLE IV: Control performance comparison on a 2-DoF planar robot.
Controller Static Dynamic
Success Rate Tra. Error Success Rate Tra. Error
PD $0.874$ $0.068\pm 0.017$ $0.112$ $0.062\pm 0.027$
PD + CBF-QP [7] $0.982$ $0.131\pm 0.049$ $0.638$ $0.217\pm 0.106$
PD + DR-CBF-QP $\boldsymbol{1.0}$ $0.173\pm 0.074$ $\boldsymbol{0.992}$ $0.394\pm 0.172$
Next, we present our evaluation of the proposed DR-CBF-QP formulation for safe control synthesis on a planar robot.
Setup: The controllers are evaluated in both static (Fig. 1) and dynamic environments (Fig. 2). In both settings, point cloud data of the obstacle surfaces is provided to the robot. For dynamic environments, obstacle velocities are sampled from a normal distribution, $v\sim\mathcal{N}(0.5,0.1)$ m/s, while the robot is provided with a nominal velocity of $v=0.5$ m/s for control synthesis. Each controller is evaluated for $500$ random trials in both static and dynamic environments.
In both scenarios, the objective is to track the bubble-CDF planned trajectory $\gamma$ (Fig. 1) while avoiding collisions.
Results and Discussion: Table IV presents the control performance across static and dynamic environments. The proposed DR-CBF-QP formulation achieves a $100\%$ success rate in static environments and high success rate ($99.2\%$) in dynamic environments, outperforming both the PD controller and the standard CBF-QP approach. The PD controller, designed solely for trajectory tracking without obstacle avoidance, exhibits the lowest success rates in both static and dynamic environments. The baseline CBF-QP shows good performance in static scenarios, but its success rate significantly deteriorates in dynamic environments ($63.8\%$). We attribute this degradation to disregarding uncertainties in dynamic obstacle velocities.
While the DR-CBF-QP formulation demonstrates higher tracking errors compared to the baselines, this trade-off reflects its emphasis on maintaining safety. The method adjusts the robot’s trajectory to prioritize collision avoidance, especially in dynamic environments where safety margins must adapt to uncertainties in obstacle motion and CDF estimates.
Qualitatively, the defensive maneuvers and adaptability of the DR-CBF-QP formulation are evident in Fig. 2. In dynamic environments, the robot effectively adjusts its trajectory to avoid collisions, as shown in Fig. 2(b), where it performs a defensive maneuver to bypass a moving obstacle. Subsequently, it resumes tracking the planned trajectory (Fig. 2(c)) and successfully reaches the goal configuration (Fig. 2(d)).
VIII-C 6-DoF xArm Robot Simulation
Next, we present simulated experiments for a 6-DoF xArm robotic manipulator in Pybullet [8] with static and dynamic obstacles to further evaluate our proposed approach.
Setup: The simulation environment consists of a 6-DoF xArm robot positioned on a table next to a shelf, as illustrated in Fig. 3. A depth camera provides point-cloud observations of the obstacles. We conducted 50 randomized trials by varying the shelf position and selecting different random goal positions. For each trial, inverse kinematics was used to determine five feasible goal configurations, and motion planning was performed in a static environment using these configurations.
To evaluate controller performance, the best planned path from the bubble-CDF planner was selected for tracking in each trial. Additionally, three dynamic obstacles were introduced into the scene, with their positions and velocities varying but assumed to be known to the robot. The velocities of the dynamic obstacles were sampled from a normal distribution, $v\sim\mathcal{N}(0.2,0.05)$ m/s.
TABLE V: Planning performance comparison on a 6-DoF xArm robot in Pybullet Simulation.
Planner Col. Checks Path Length
CDF-RRT $3415.4\pm 803.6$ $3.11\pm 0.46$
SDF-RRT [46] $3251.8\pm 782.4$ $3.07\pm 0.49$
SDF-RRT-Connect $3437.4\pm 723.2$ $3.11\pm 0.43$
SDF-Lazy-RRT $2741.1\pm 737.9$ $3.15\pm 0.53$
Bubble-CDF $\boldsymbol{278.5\pm 80.7}$ $3.04\pm 0.46$
(a) Initial Configuration
(b) Planned Configuration 1
(c) Planned Configuration 2
(d) Goal Configuration
Figure 3: Bubble-CDF planning for a 6-DoF xArm robot in a static environment, targeting an end-effector goal represented by a green sphere. (a) The initial configuration of the xArm. (b, c) Intermediate configurations illustrating the planned path as the robot avoids obstacles while progressing toward the goal. (d) The final goal configuration reached by the robot.
(a) Start of Execution
(b) Dynamic Obstacle Approaching
(c) Defensive Maneuver
(d) Goal Reached
Figure 4: Snapshots of safe control execution on a 6-DoF xArm robot in an environment with dynamic obstacles. (a) The control execution begins at the initial configuration. (b) A dynamic obstacle (blue) approaches the robot from right. (c) The robot executes a defensive maneuver, moving upward to avoid the obstacle. (d) The robot successfully resumes tracking and reaches the goal configuration.
Results and Discussion: Table V summarizes the planning results for the 6-DoF xArm robot. Similar to the 2-DoF case, the bubble-CDF planner significantly outperforms baseline methods in terms of collision checks. Path lengths across all planners remain comparable, indicating that the bubble-CDF approach does not compromise solution quality.
Figure 3 illustrates the planned path generated by the bubble-CDF planner in a static environment for a specific end-effector goal inside the shelf. Among the five possible goal configurations derived from inverse kinematics, the planner selects the shortest collision-free path.
Table VI quantitatively compares the control performance of different controllers under both static and dynamic environments. The baseline PD controller achieves a success rate of 68% in the static environment but struggles significantly in dynamic scenarios, with a success rate dropping to 18%, highlighting its inability to effectively react to moving obstacles.
The addition of a CBF-QP safety filter improves the success rate to 88% in static environments. However, in dynamic scenarios, its success rate remains limited at 70%, suggesting that while CBF constraints enhance obstacle avoidance, they do not account for errors in CDF modeling and point-cloud observations. In contrast, our proposed PD + DR-CBF-QP achieves a 100% success rate in both static and dynamic environments, demonstrating its robustness against dynamic obstacles and uncertainty. The increased tracking error reflects the adaptive nature of our approach, which prioritizes safety by dynamically adjusting the trajectory to avoid obstacles, even if it deviates from the originally planned path.
Figure 4 illustrates the execution of a planned path by the 6-DoF xArm robot in a dynamic environment. The robot begins at the initial configuration (Fig. 4(a)) and dynamically reacts to an approaching obstacle (blue) as shown in Fig. 4(b). By executing a defensive maneuver (Fig. 4(c)), the robot moves upward to maintain safety and avoid a collision. After the obstacle passes, the robot resumes trajectory tracking and successfully reaches the goal configuration (Fig. 4(d)).
TABLE VI: Control performance comparison on a 6-DoF xArm robot in Pybullet Simulation.
Controller Static Dynamic
Success Rate Tra. Error Success Rate Tra. Error
PD $0.68$ $0.053\pm 0.014$ $0.18$ $0.051\pm 0.013$
PD + CBF-QP [7] $0.88$ $0.115\pm 0.046$ $0.70$ $0.327\pm 0.104$
PD + DR-CBF-QP $1.0$ $0.143\pm 0.055$ $1.0$ $0.516\pm 0.221$
VIII-D 6-DoF xArm Robot Experiments
TABLE VII: Planning performance comparison on a real 6-DoF xArm robot.
Planner Col. Checks Path Length
CDF-RRT $5617.4\pm 1333.6$ $3.71\pm 0.71$
SDF-RRT [46] $5911.9\pm 1487.7$ $3.72\pm 0.68$
SDF-RRT-Connect $5801.5\pm 1265.4$ $3.75\pm 0.63$
SDF-Lazy-RRT $4765.8\pm 1331.4$ $3.84\pm 0.69$
Bubble-CDF $\boldsymbol{634.8\pm 225.1}$ $3.81\pm 0.66$
(a) Planned configurations
(b) Response to dynamic obstacle
(c) Response to dynamic obstacle
(d) Response to dynamic obstacle
(e) Planned configurations
(f) Response to dynamic obstacle
(g) Response to dynamic obstacle
(h) Response to dynamic obstacle
Figure 5: Bubble-CDF planner and DR-CBF control applied to two real-world setups for a 6-DoF xArm robot. The top row (a-d) represents Setup 1, with the robot navigating a cluttered environment featuring a combination of static and dynamic obstacles. Similarly, the bottom row (e-h) depicts Setup 2, showcasing the planner’s adaptability in a different obstacle layout. For both setups: (a, e) illustrate the bubble-CDF planned configuration in static environments, while (b-d, f-h) demonstrate the robot’s real-time adaptive responses to dynamic obstacles.
To further validate the efficiency of the bubble-CDF planner and the robustness of the DR-CBF controller, we conducted real-world experiments on a 6-DoF xArm manipulator in several cluttered and dynamic environments.
Setup: The experimental environment consists of a 6-DoF xArm robot operating on a table in a cluttered workspace with static obstacles, as shown in Fig. 5. A depth camera provides point-cloud observations of the scene. We conducted 20 randomized trials by varying obstacle placements and selecting different goal configurations. Planning was performed assuming the static environment. During execution, the robot followed the planned path while dynamic obstacles were introduced. The velocity of the moving objects is estimated using ArUco markers.
Results and Discussion: Table VII presents a quantitative comparison of planning performance across different methods. The bubble-CDF planner significantly reduces the number of collision checks compared to all baselines, achieving a nearly tenfold reduction. This improvement stems from the use of configuration-space bubbles, which certify local collision-free regions and minimize the need for frequent edge-based collision checking. Consistent with simulation results, the generated path lengths remain comparable across all methods, demonstrating that our approach maintains solution quality while substantially improving efficiency.
Figure 5 shows snapshots of the bubble-CDF planner and DR-CBF controller in two different workspace arrangements. The robot exhibits similar behavior to the simulation, following the nominal planned trajectory (Figs. 5(a)5(e)) but stopping and avoiding dynamic obstacles as they are presented. Video recordings of these experiments are shown in the supplementary material.
IX Limitations and Future Directions
While our proposed bubble-CDF planner and DR-CBF safety filter demonstrate strong performance in motion planning and control for robotic manipulators, several limitations demand further investigation. First, a fundamental challenge lies in balancing safety and stability, particularly in highly cluttered environments. While the DR-CBF-QP formulation ensures collision-free execution, it does not explicitly account for cases where the robot may become stuck in front of obstacles, unable to progress toward the planned path. This incompatibility between obstacle avoidance and goal-reaching poses a significant challenge, and future work should focus on addressing the compatibility of safety and stability.
Second, although our formulation is designed to handle uncertainty in CDF modeling, point-cloud observations, and dynamic obstacle velocity estimates, reliable obstacle point clouds and velocity estimation remains important for real-world deployments. In real-world experiments, we currently rely on ArUco markers for velocity tracking, but a more general and robust approach, such as learning-based perception or multi-frame point-cloud tracking, could improve adaptability in unstructured environments.
Third, the current framework does not explicitly consider self-collision, which may lead to infeasible or unsafe configurations where the robot intersects with itself. Incorporating self-collision constraints into the neural CDF representation or integrating explicit collision-checking mechanisms within the motion planner would enhance the framework’s applicability to more complex robotic systems.
Finally, the use of configuration-space bubbles for planning, while computationally efficient, may be overly conservative in densely cluttered environments. Since spherical safe regions do not fully exploit free space, alternative representations such as ellipsoidal or polyhedral safe regions could enable more flexible and efficient motion planning.
X Conclusion
In this work, we proposed an integrated framework for safe motion planning and control of robotic manipulators in dynamic and cluttered environments. Our approach combines a novel bubble-CDF planner for efficient trajectory generation and a distributionally robust control barrier function (DR-CBF) formulation for safe navigation. Comprehensive evaluations, conducted in both simulation and real-world settings, validate the effectiveness of our approach, demonstrating high success rates and significant reductions in computational overhead.
The bubble-CDF planner leverages the neural configuration-space distance function (CDF) to construct local safety regions, termed configuration-space bubbles, enabling efficient exploration of the configuration space. By certifying entire regions as collision-free with a single query, the planner significantly reduces the number of collision checks compared to conventional sampling-based methods. This efficiency is achieved without compromising the quality of the planned trajectories, as evidenced by comparable path lengths to baseline approaches. Additionally, the planner allows trajectory optimization to be formulated as a convex program, enabling smooth and safe trajectory refinement without requiring additional collision checks.
For real-time trajectory execution, we introduced a DR-CBF-QP safety filter that ensures the robot remains within a safe set throughout its motion. This formulation explicitly incorporates uncertainties in the neural CDF model, gradient estimates, and sensor measurements, leveraging a distributionally robust optimization framework. The DR-CBF-QP controller dynamically adapts to obstacles and sensor inaccuracies, maintaining safety while ensuring smooth trajectory tracking in both static and dynamic environments.